OpenStudy (anonymous):

The container that holds water for the football team is 1/2 full. After pouring in 6 gallons of water, it is 7/10 full. How many gallons of water can the container hold?

OpenStudy (campbell_st):

well if you look at it 6 gallons = 7/10 - 5/10 so 6 gallons = 2/10 of the total... perhaps multiplying by 5 may give the solution...?

OpenStudy (anonymous):

Thanks for the help.